OpenStudy (anonymous):

A business has $11,080 to spend on new laptops and tablet computers for its salespeople. The laptops cost $515 each. The tablets cost $285 each. The business wants each salesperson to have either a laptop or a tablet. There are 30 salespeople. How many of each type of computer should the business buy? 9 laptops and 21 tablets 11 laptops and 19 tablets 19 laptops and 11 tablets 21 laptops and 9 tablets

OpenStudy (beccaboo333):

Give me a moment.

OpenStudy (beccaboo333):

not the first option

OpenStudy (beccaboo333):

second

OpenStudy (beccaboo333):

because 11x515=5665 19x285=5415 5415+5665=11080

OpenStudy (anonymous):

515L + 285T = 11080 L + T = 30 ---> L = 30 _ T 515(30 - T) + 285T = 11080 15450 - 515T + 285T = 11080 -515T + 285T = 11080 - 15450 -230T = - 4370 T = -4370/-230 T = 19 L + T = 30 L + 19 = 30 L = 30 - 19 L = 11 19 tablets and 11 laptops

OpenStudy (beccaboo333):

O.o

OpenStudy (anonymous):

oops...L = 30 - T

OpenStudy (jdoe0001):

\(\large { \color{seagreen}{ l+t=30} \Leftarrow \textit{there are 30 salespersons}\\ \quad \\ 515\times l+285\times t=11,080\Leftarrow \begin{array}{llll} \textit{laptops are 515}\\ \textit{tablets are 285} \end{array}\\ \quad \\ \color{seagreen}{ 515l+285t=11,080}}\) solve the system of equation for either variable